OpenStudy (anonymous):

Consider this system of linear equations: 22x − y = -23 (equation 1) 11x + 12y = 1 (equation 2) To find the value of x, you need to substitute in equation 2. The value of x is .

OpenStudy (kidrah69):

ok well fist you need to get X equals or a y equals equation

OpenStudy (kidrah69):

Do you know how to do that?

OpenStudy (anonymous):

no

OpenStudy (kidrah69):

Hint: since the first one has a varible only and no number nxt to it it will be easier

OpenStudy (kidrah69):

\[-y=-22x-23\] here i did that step now just divide everything by -1 to get a positive Y value

OpenStudy (anonymous):

how would we get a positive out of this

OpenStudy (kidrah69):

\[\frac{ −y=−22x−23 }{ -1=-1-1~~ }\]

OpenStudy (kidrah69):

A negative times a negative is a positive same works for division

OpenStudy (anonymous):

wouldn't it be 22x and 23 because -1 can go into -22x 22 time right?

OpenStudy (kidrah69):

nope the equation stays the same the signs flip :P

OpenStudy (anonymous):

i dont understand

OpenStudy (kidrah69):

\[y=22x+23\]

OpenStudy (kidrah69):

thats all

OpenStudy (anonymous):

okay but what about the value of x?

OpenStudy (kidrah69):

Well we have out equation ONE so now substitute! \[11x + 12(22x+23) = 1 \] this will get you the X value Solve.

OpenStudy (anonymous):

wouldn't it be 1

OpenStudy (kidrah69):

now 1 but -1

OpenStudy (anonymous):

oh so -1 is my answer

OpenStudy (kidrah69):

now all you do is plug the X value in equation 1 \[y=22(-1)+23\] yes the x=-1